OpenStudy (anonymous):

An LCO Flat Screen TV at Best Buy increased in price from $900 to $1,200. What was the percent of increase

OpenStudy (anonymous):

33 1/3%

OpenStudy (anonymous):

how did you get to that answer?

OpenStudy (anonymous):

% increase = (1,200 - 900)/ 900 = 3/9 = 1/3 = 33%

OpenStudy (anonymous):

thanks!